OpenStudy (anonymous):

Solve the following system of linear equations using either the substitution method or the addition method: A.) 2x+y=5 B.) 4x+3y=14

OpenStudy (solomonzelman):

What do you choose substitution or addition?

OpenStudy (solomonzelman):

I would do addition, multiply the first equation by '-2'. 2x+y=5 ---> 2x(-2)+y(-2)=5(-2) ---> -4x-2y-10 4x + 3y =14 -4x-2y=-10 add them to each other.

OpenStudy (anonymous):

addition yeah is what I was looking to do